OpenStudy (anonymous):

Find the solution of each of the following differential equations for a function x = x(t): dx/dt − 2xt = t^2 − 1 (Hint: Find an integrating factor) Your assistance is appreciated. Kind regards

OpenStudy (anonymous):

So far: I rewrite the equation as : dx/dt = -2/t(x) = t^2-1

OpenStudy (anonymous):

integrating factor , using -2/t is : e^integral(-2/t dt

OpenStudy (anonymous):

e^integral(-2/t dt = e^-2lnt = -2t

OpenStudy (anonymous):

Is this correct so far?

OpenStudy (anonymous):

You should ask this question in the math group. Do you know how to access the math group?

OpenStudy (anonymous):

Oh dear, can I move it there..lol

OpenStudy (anonymous):

Nope. You'll have to close it here & re-type it there..lol.

OpenStudy (anonymous):

Thank you...I was wondering why no replies...hehehe

OpenStudy (anonymous):

lol.